OpenStudy (paradise):

What is the seventh term of (x + 4)^8

OpenStudy (anonymous):

(a + b)^n = sum(k=0 to n) C(n, k) * a^(n - k) * b^k. So we have: (x + 4)^8 = sum(k=0 to 8) C(8, k) * x^(8 - k) * 4^k. Then, we see that the seventh term will occur at k = 6. Therefore, the seventh term is C(8, 7) * x^(8 - 7) * 4^6 = 32768x^2.

OpenStudy (anonymous):

does that help??

OpenStudy (paradise):

that's the same answer I keep getting but its not in one of my choices

OpenStudy (anonymous):

this is easy

OpenStudy (anonymous):

it is 34

OpenStudy (paradise):

A. 114,688x2 B. 114,688x3 C. 114,688x4 D. 114,688x5 Its one of these

OpenStudy (anonymous):

but the whole point of this website is to get the answer by working out the problem but okayyy :/

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

what math are u in

OpenStudy (anonymous):

dats what i was gonna ask lol^

OpenStudy (anonymous):

lol

OpenStudy (paradise):

Algebra 2 haha

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

cant help u there

OpenStudy (anonymous):

well for me im startn algebra 2 but im somewhat good at it when i comes to certain parts but im really confused!!! :l

OpenStudy (paradise):

Ya same here /:

OpenStudy (anonymous):

dont follow me follow what u knoe im think a but at da same time itz prolow b cuz im thinking simple how many time can4 go into 8? just thinking :/

OpenStudy (anonymous):

best guess go wit c lol

OpenStudy (anonymous):

@ChiefSosa3Hunna u on hush or u waiting for her reply

OpenStudy (anonymous):

itz a... im smart!!!

OpenStudy (anonymous):

im just sittn her sippn on some lemonade :)

OpenStudy (paradise):

haha gotta love lemonade

OpenStudy (anonymous):

when u got lemons made lemonade or lemon juice :/

OpenStudy (anonymous):

ok r u doing a binomial theorem assesment ?

OpenStudy (paradise):

umm its Operations with Polynomials

OpenStudy (anonymous):

translation he smokkin on tht kush lol

OpenStudy (paradise):

hahaha!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol wow i dont smoke well i stopd for a day

OpenStudy (anonymous):

lol surez

OpenStudy (anonymous):

well I gtg see yall later

OpenStudy (anonymous):

byeee!!!!!! :O

OpenStudy (paradise):

haha for a day